# Minimal multiplicity of fiber components in abelian fibrations

### v1

@@ 1 Introduction
On the other hand, examples of fibrations with fibers higher genus curves for which the two versions differ are found in [W], [S], [Ca3]. A remarkable example is given in [L], where the smooth fibers are two Enriques surfaces of multiplicities 2 and 3, with consequences in hyperbolic and arithmetic geometries derived in [BCJW].
@@
Theorem 1.2: (3) Let $f:M\to B$ be a proper Kähler fibration with $M$ smooth, $B$ normal, such that its generic fibers are smooth complex tori. Assume that $f$ has equidimensional fibers. Then for any $b\in B$, $f$ has a meromorphic section above some open neighborhood of $b$ in $B$.
@@ 3 Higher dimensional bases
Theorem 3.1: Let $f:M\to B$ be a proper Kähler fibration between normal complex spaces such that its generic fibers are smooth compact complex tori. Assume that $f$ has equidimensional fibers. Then for any $b\in B$, $f$ has a meromorphic section above some open neighborhood of $b$ in $B$.
@@
Recall ([BM], chap IV, Théorème 3.4.1, Corollaire 3.4.2) that $f$ being equidimensional, with $B$ normal, the components of the fibers of $f$ can be given in a unique way multiplicities (equal to $1$ for those above $B^{0}$) in such a way that the fibers of $f$ form an analytic family of cycles in the Barlet sense, such that the map $\beta:B\to C(M/B)$ sending $M_{b}$ to its reprensentative point in $C(M/B)$, the Chow/Barlet space of relative $f$-cycles of $M$, is analytic. Moreover, if $C:=\beta(B)\subset C(M/B)$, and if $\Gamma_{C}\subset C\times M$ is the incidence graph of the universal family of cycles parametrised by $C$, then $(f\times id_{X})(M)=(\beta\times id_{X})^{-1}(\Gamma_{C})$: by its very definition, the graph of $f$ is obtained from $\Gamma_{C}$ by the base-change $\beta$.
Let now $b^{\prime}\in B\setminus B^{0}$, and $M_{b^{\prime}}=\sum_{i}m_{i}\cdot F_{i}$ be its $f$-fiber. By normality of $M$, through the generic point of each $F_{i}$ we have a local multisection $s_{i},i=1,\dots,r$ of $f$ which has degree $m_{i}$ over $B$. Denote with $p_{i,j}(b)$ its points over $b$. By hypothesis, the gcd of the numbers $\{m_{i}\}$ is one, and Bezout lemma implies the existence of integers $v_{i}$ such that $\sum_{i}u_{i}:=\sum_{i}v_{i}\cdot m_{i}=1,u_{i}:=v_{i}\cdot m_{i}$. By averaging (4.2), for each $b\in B^{0}$ near $b^{\prime}$, a well-defined point $s(b):=\sum_{i,j}u_{i}\cdot p_{i,j}(b)\in M_{b}$, which depends holomorphically on $b$.
We shall show, using the Kählerianity assumption, that this section of $f$ over $B^{0}$ near $b^{\prime}$ extends meromorphically to $B$ near $b^{\prime}$. More precisely, we will show that the addition (and subtraction) in the generic smooth fibers of $f$ extend meromorphically to $M\times_{B}M$. Any section obtained by averaging over the locus $B^{0}$ of smooth fibers of $f$ in $B$ thus extends meromorphically to $B$.
@@
2. The hypothesis on the $\operatorname{\text{\it gcd}}$ of the $m_{i}$ being equal to $1$ do not follow from the same property in codimension $1$, even if $M$ and $B$ are smooth, $f$ projective, and if the fibers outside $b\in B$ are reduced, as shown by the example of I. Hellmann ([Hel, Section 2.2]).
@@ 4.2 The averaging formalism
Acknowledgments. L. K. and M. V. are grateful to J. Kollár for his insightful comments and email exchanges, and to F. Bogomolov, whose ideas were of great influence in this paper.

### v2

@@ 1 Introduction
On the other hand, examples of fibrations with fibers higher genus curves for which the two versions differ are found in [W], [S], [Ca3]. A remarkable example is given in [La], where the smooth fibers are Enriques surfaces, but some singular fibre has components of multiplicities 2 and 3, the base being the projective line, with consequences in hyperbolic and arithmetic geometries derived in [BCJW].
@@
When $\dim_{\mathbb{C}}M=2$, 2 is a consequence of Kodaira’s classification of elliptic fibrations, [BHPV]. After we posted this paper on the arXiv, we became aware that 2 was already proved in an unpublished paper of S. Lu, [Lu].
@@
Theorem 1.2: (3) Let $f:M\to B$ be a proper Kähler fibration with $M$ smooth, $B$ normal, such that its generic fibers are smooth complex tori. Assume that $f$ has equidimensional fibers. Let $b\in B$ be any point and let $F_{1},...,F_{k}$ be the irreducible components of $f^{-1}(b)$, and $\mu_{1},...,\mu_{k}$ their multiplicities. Assume that the greatest common divisor of $\mu_{i}$ is 1. Then for any $b\in B$, $f$ has a meromorphic section above some open neighborhood of $b$ in $B$.
@@ 3 Higher dimensional bases
Theorem 3.1: Let $f:M\to B$ be a proper Kähler fibration between normal complex spaces, $M$ smooth, such that its generic fibers are smooth compact complex tori. Assume that $f$ has equidimensional fibers. Let $b\in B$ be any point and let $F_{1},...,F_{k}$ be the irreducible components of $f^{-1}(b)$, and $\mu_{1},...,\mu_{k}$ their multiplicities. Assume that the greatest common divisor of $\mu_{i}$ is 1. Then $f$ has a meromorphic section above some open neighborhood of $b$ in $B$.
@@
Recall ([BM], chap IV, Théorème 3.4.1, Corollaire 3.4.2) that $f$ being equidimensional, with $B$ normal, the components of the fibers of $f$ can be given in a unique way multiplicities (equal to $1$ for those above $B^{0}$) in such a way that the fibers of $f$ form an analytic family of cycles in the Barlet sense, such that the map $\beta:B\to C(M/B)$ sending $M_{b}$ to its representative point in $C(M/B)$, the Chow/Barlet space of relative $f$-cycles of $M$, is analytic. Moreover, if $C:=\beta(B)\subset C(M/B)$, and if $\Gamma_{C}\subset C\times M$ is the incidence graph of the universal family of cycles parametrised by $C$, then $(f\times id_{X})(M)=(\beta\times id_{X})^{-1}(\Gamma_{C})$: by its very definition, the graph of $f$ is obtained from $\Gamma_{C}$ by the base-change $\beta$.
Let now $b^{\prime}\in B\setminus B^{0}$, and $M_{b^{\prime}}=\sum_{i}m_{i}\cdot F_{i}$ be its $f$-fiber. By smoothness of $M$, through the generic point of each $F_{i}$ we have a local multisection $s_{i},i=1,\dots,r$ of $f$ which has degree $m_{i}$ over $B$. Denote with $p_{i,j}(b)$ its points over $b$. By hypothesis, the gcd of the numbers $\{m_{i}\}$ is one, and Bezout lemma implies the existence of integers $v_{i}$ such that $\sum_{i}u_{i}:=\sum_{i}v_{i}\cdot m_{i}=1,u_{i}:=v_{i}\cdot m_{i}$. By averaging (4.2), for each $b\in B^{0}$ near $b^{\prime}$, a well-defined unique point $s(b):=\sum_{i,j}u_{i}\cdot p_{i,j}(b)\in M_{b}$ exists, which depends holomorphically on $b$.
We shall show, using the Kählerianity assumption, that this section of $f$ over $B^{0}$ near $b^{\prime}$ extends meromorphically to $B$ near $b^{\prime}$. More precisely, we will show that the addition (and substraction) in the generic smooth fibers of $f$ extend meromorphically to $M\times_{B}M$. Any section obtained by averaging over the locus $B^{0}$ of smooth fibers of $f$ in $B$ thus extends meromorphically to $B$.
@@
2. The hypothesis on the $\operatorname{\text{\it gcd}}$ of the $m_{i}$ being equal to $1$ does not follow from the same property in codimension $1$, even if $M$ and $B$ are smooth, $f$ projective, and if the fibers outside $b\in B$ are reduced, as shown by the example of I. Hellmann ([Hel, Section 2.2]).
@@ 4.2 The averaging formalism
Acknowledgments. L. K. and M. V. are grateful to J. Kollár for his insightful comments and email exchanges, and to F. Bogomolov, whose ideas were of great influence in this paper. We express our gratitude to J. Kollàr for noticing that, in the first version of this paper, the gcd hypothesis had been forgotten in the statement of Theorem 3. We are also grateful to S. Lu for communicating to us the results of [Lu].